14) how to solve 3x+y=-8 with the substitution of -2x-y=6
Answer by stanbon(75887)

how to solve
3x+y=-8
with the substitution of
-2x-y=6
----
Add the two equations and solve for "x":
x = -2
============
Solve for "y":
3x+y = -8
3(-2)+y = -8
-6 + y = -8
y = -2
==========
Cheers,
Stan H.